# Phase-resolving the absorption signatures of water and carbon monoxide in the atmosphere of the ultra-hot Jupiter WASP-121b with GEMINI-S/IGRINS

### v1

@@ 1 Introduction
The first atmospheric studies of WASP-121b were performed with HST/WFC3, at low spectral resolution. Evans et al. (2016) reported water in the planet’s transmission spectrum, while Evans et al. (2017) found emission features associated with water in its secondary-eclipse spectrum, indicating a thermal inversion on the dayside (e.g., Hubeny et al. 2003; Fortney 2005; Gandhi & Madhusudhan 2019). They also reported evidence for VO, an optical absorber potentially contributing towards this inversion, but its presence in emission was later ruled out by additional observations with the same instrument (Mikal-Evans et al. 2019, 2020). Conversely, optical transmission spectra taken with HST/STIS and WFC3 (Evans et al. 2018) did lead to constraints on the VO abundance, possibly because these probe a different part of the atmosphere. The data also revealed strong excess absorption at near-UV wavelengths. At the same time, no evidence for TiO (another frequently postulated inversion-agent) was found, indicating that Ti-bearing species may be cold-trapped on the nightside of WASP-121b (e.g., Spiegel et al. 2009; Parmentier et al. 2013; Hoeijmakers et al. 2022).
In recent years, ground-based high-resolution spectroscopy (HRS; 15,000) has painted a more detailed picture of the chemical inventory of WASP-121b. To date, HRS has led to the unambiguous detection of Na i (Sindel et al. 2018), Fe i (Gibson et al. 2020), H i (Cabot et al. 2020), Cr i, V i, Fe ii (Ben-Yami et al. 2020), Mg i, Ca i, Ni i (Hoeijmakers et al. 2020), K i, Li i, Ca ii (Borsa et al. 2021), Sc ii (Merritt et al. 2021), Mn i (Hoeijmakers et al. 2022), Co i, Sr ii, Ba ii (Azevedo Silva et al. 2022), CO, and H2O (this work)111Of the studies listed here, only the observations by Hoeijmakers et al. (2022) were in emission. . Notably, none of these studies detected Ti i, TiO or VO in the transmission or emission spectrum of WASP-121b (see also Merritt et al. 2020). While the non-detection of Ti i and TiO is consistent with the cold-trapping hypothesis, Hoeijmakers et al. (2020, 2022) argued that their detection of V i should imply the presence of VO, but that their analyses were not sensitive enough to the latter. A more recent detection of V i and VO in the transmission spectrum of the canonical UHJ WASP-76b (Pelletier et al. 2023) supports the idea that these species exist together.
@@
2358 52 K | D16 |
Phase-curve observations with TESS (Bourrier et al. 2020b; Daylan et al. 2021), HST/WFC3 (Mikal-Evans et al. 2022), Spitzer/IRAC (Morello et al. 2023), and, most recently, JWST/NIRSpec (Mikal-Evans et al. 2023) have yielded important constraints on the global temperature structure of WASP-121b, and the contrast between the dayside and the nigthside. While the (optical) TESS phase curves are consistent with a hotspot at the substellar point, the (infrared) HST and JWST phase curves show a peak just before secondary eclipse (3∘ in phase for JWST), suggesting a small eastward hotspot offset (assuming no clouds). Remarkably, the two Spitzer phase curves presented in Morello et al. (2023) indicate a westward offset, possibly due to instrument systematics. An independent re-analysis of the same data by Davenport et al. (subm.) reports offsets that are consistent with zero and close to the JWST results (but less precise). At the same time, there is a growing body of literature claiming signs of temporal variability in the atmosphere of WASP-121b (Wilson et al. 2021; Ouyang et al. 2023; Changeat et al. 2024), which could be due to weather cycles on the planet.
@@ 2.3 Correcting for telluric absorption
There is no golden rule prescribing how many SVD components to remove from the data – removing too few will cause systematics to persist, while removing too many will affect the planet signal. Moreover, recent work by Smith et al. (2024) on IGRINS data shows that it may be advantageous to remove a different number of SVD components per order and per night, depending on the observing conditions (humidity, airmass, etc.) and the degree of telluric absorption within an order555For further discussion, see Nugroho et al. (2017), Spring et al. (2022), Cheverall et al. (2023), and Klein et al. (2024).. However, designing such an optimisation process is non-trivial, and HRS studies generally opt to remove the same number of SVD components from each order for a given observing night (e.g., Line et al. 2021; Van Sluijs et al. 2022; Boucher et al. 2023; Brogi et al. 2023; Pelletier et al. 2023; Smith et al. 2024). In this work, we adopt a similar approach. Yet, to remain as agnostic as possible with regards to the telluric correction, we carry out separate analyses for , with the fixed number of SVD components removed from each of the orders, for each night. The motivation for the lower limit of = 3 is that, from this value onwards, the – maps of the combined transits (see Fig. 3) show clear detections of CO and HO near the expected planet position. The upper limit = 18 was chosen because this is where the SNR associated with the HO detection first drops below 4 (Fig. 4).
@@ 5.1 Four models of WASP-121b
To explore the effect of atmospheric drag (e.g., Showman et al. 2013; Komacek & Showman 2016; Parmentier & Crossfield 2018), we also consider two models with drag timescales s (weak drag) and s (strong drag), which are shown in the third and fourth column of Fig. 8. The drag timescale encapsulates a variety of physical mechanisms, such as turbulent mixing (Li & Goodman 2010), Lorentz-force braking of ionised winds in the planet’s magnetic field (Perna et al. 2010a), and Ohmic dissipation (Perna et al. 2010b). The drag deposits energy back into the atmosphere. As shown in Fig. 8, increasing the drag strength (and thus lowering ) slows down winds in the atmosphere131313 is a tunable parameter in the GCM. It is included as an additional “Rayleigh drag” term in the momentum equation solved by the dynamical core. This approach assumes that the drag force is uniform across the whole planet atmosphere. A drag-free scenario corresponds to .. This leads to less efficient heat redistribution and a more symmetric temperature structure. Also, the equatorial jet gets suppressed – in the strong drag-model, there is only a day-to-night flow.
Another notable difference between the model from Parmentier et al. (2018) and the models from Tan et al. (2024) is that the latter account for opacities from Fe when evaluating heating and cooling rates, making the thermal inversion on the dayside extend to much lower pressures. We refer to Table 1 in Tan et al. (2024) for the full list of opacities considered in their radiative transfer and to Freedman et al. (2014) for the opacities used by Parmentier et al. (2018). While the extra optical opacity this causes the models with H dissociation/recombination to look substantially hotter, they have nearly the same effective temperature (2400 K) at the dayside photosphere (0.1 bar) as the model from Parmentier et al. (2018).
@@ 5.2 Computing absorption trails
Before feeding the GCM outputs into gCMCRT, we map the atmospheric structures onto a 3D grid with altitude (instead of pressure) as a vertical coordinate. We account for the fact that each atmospheric column has a different scale height, which is set by local gravity, temperature, and mean-molecular weight. For each of the four WASP-121b models we simulate 25 spectra (equidistant in orbital phase) between phase angles 17.6, which also cover the ingress and egress. We assume an edge-on orbit, a semi-major axis of 0.025 AU, a stellar radius of 1.46 , and an orbital period of 1.27 days. At each orbital phase angle, gCMCRT simulates a transmission spectrum by randomly shooting photon packets at the planet limb and evaluating the optical depth encountered by each photon packet. In this calculation, it accounts for the Doppler shifts imparted on the opacities by the radial component of the local wind vector and planet rotation (Wardenier et al. 2021). The transit depth at a given wavelength is then obtained by averaging over all photon packets. We use 105 photon packets per wavelength to accurately model the shapes, depths, and shifts of the spectral lines. Because we do not explicitly treat scattering, the propagation direction of the photon packets does not change throughout the calculation. As in Wardenier et al. (2023), the optical and infrared spectra are computed at resolutions and , respectively. For the radiative transfer, we include the same set of continuum opacities and line species as in Wardenier et al. (2023).
@@ 6.6 Non-detection of OH
Although we find a hint of OH in the atmosphere of WASP-121b (see Figs. 3, 14), the associated signal (SNR 2.5) is far from significant. This is remarkable given that OH was confidently detected (SNR 6.1) in the atmosphere of WASP-76b based on one transit visit with CARMENES (; Landman et al. 2021), which is on a smaller, 3.5-m telescope. Since the equilibrium temperature of WASP-121b is 250 K higher than that of WASP-76b, one explanation could be that OH has further dissociated into atomic O and H on the hotter dayside of WASP-121b, such that its features in the transmission spectrum are weaker. A retrieval would allow to shed light on the upper limits of the OH abundance in the region probed by our transit data.
@@ Appendix A Supplementary figures
Figure 15 shows the results of the VULCAN simulation described in Section 6.3. As demonstrated in the plot, chemical kinetics only has a strong impact on the water abundances in small regions around 10-5 bar, which is not enough to substantially impact the HO signal of the planet.

### v3

@@ 1 Introduction
The first atmospheric studies of WASP-121b were performed with HST/WFC3, at low spectral resolution. Evans et al. (2016) reported water in the planet’s transmission spectrum, while Evans et al. (2017) found emission features associated with water in its secondary-eclipse spectrum, indicating a thermal inversion on the dayside (e.g., Hubeny et al. 2003; Fortney 2005; Gandhi & Madhusudhan 2019). They also reported evidence for VO, an optical absorber potentially contributing towards this inversion, but its presence in emission was later ruled out by additional observations with the same instrument (Mikal-Evans et al. 2019, 2020). Conversely, optical transmission spectra taken with HST/STIS and WFC3 (Evans et al. 2018) did lead to constraints on the VO abundance, probably because these observations probe a different part of the atmosphere. The data also revealed strong excess absorption at near-UV wavelengths. At the same time, no evidence for TiO (another frequently postulated inversion-agent) was found, indicating that Ti-bearing species may be cold-trapped on the nightside of WASP-121b (e.g., Spiegel et al. 2009; Parmentier et al. 2013; Hoeijmakers et al. 2024).
In recent years, ground-based high-resolution spectroscopy (HRS; 15,000) has painted a more detailed picture of the chemical inventory of WASP-121b. To date, HRS has led to the unambiguous detection of Na i (Sindel et al. 2018), Fe i (Gibson et al. 2020), H i (Cabot et al. 2020), Cr i, V i, Fe ii (Ben-Yami et al. 2020), Mg i, Ca i, Ni i (Hoeijmakers et al. 2020), K i, Li i, Ca ii (Borsa et al. 2021), Sc ii (Merritt et al. 2021), Mn i (Hoeijmakers et al. 2024), Co i, Sr ii, Ba ii (Azevedo Silva et al. 2022), CO, and H2O (this work)111Of the studies listed here, only the observations by Hoeijmakers et al. (2024) were in emission. . Notably, none of these studies detected Ti i, TiO or VO in the transmission or emission spectrum of WASP-121b (see also Merritt et al. 2020). While the non-detection of Ti i and TiO is consistent with the cold-trapping hypothesis, Hoeijmakers et al. (2020, 2024) argued that their detection of V i should imply the presence of VO, but that their analyses were not sensitive enough to the latter. A more recent detection of V i and VO in the transmission spectrum of the canonical UHJ WASP-76b (Pelletier et al. 2023) supports the idea that these species exist together.
@@
2358 52 K | D16 | ||
217.65 km/s | B21 |
Phase-curve observations with TESS (Bourrier et al. 2020b; Daylan et al. 2021), HST/WFC3 (Mikal-Evans et al. 2022), Spitzer/IRAC (Morello et al. 2023), and, most recently, JWST/NIRSpec (Mikal-Evans et al. 2023) have yielded important constraints on the global temperature structure of WASP-121b, and the contrast between the dayside and the nigthside. While the (optical) TESS phase curves are consistent with a hotspot at the substellar point, the (infrared) HST and JWST phase curves show a peak just before secondary eclipse (3∘ in phase for JWST), suggesting a small eastward hotspot offset (assuming no clouds). Remarkably, the two Spitzer phase curves presented in Morello et al. (2023) indicate a westward offset, possibly due to instrument systematics. An independent re-analysis of the same data by Davenport et al. (subm.) reports offsets that are consistent with zero and close to the JWST results (but less precise). At the same time, there is a growing body of literature claiming signs of temporal variability in the atmosphere of WASP-121b (Wilson et al. 2021; Ouyang et al. 2023; Changeat et al. 2024). While weather cycles could play a role, we note that atmospheric variability on exoplanets is not well understood and largely remains an open question.
@@ 2.3 Correcting for telluric absorption
There is no golden rule prescribing how many SVD components to remove from the data – removing too few will cause systematics to persist, while removing too many will affect the planet signal. Moreover, recent work by Smith et al. (2024) on IGRINS data shows that it may be advantageous to remove a different number of SVD components per order and per night, depending on the observing conditions (humidity, airmass, etc.) and the degree of telluric absorption within an order555For further discussion, see Nugroho et al. (2017), Spring et al. (2022), Cheverall et al. (2023), and Klein et al. (2024).. However, designing such an optimisation process is non-trivial, and HRS studies generally opt to remove the same number of SVD components from each order for a given observing night (e.g., Line et al. 2021; Van Sluijs et al. 2023; Boucher et al. 2023; Brogi et al. 2023; Pelletier et al. 2023; Smith et al. 2024). In this work, we adopt a similar approach. Yet, to remain as agnostic as possible with regards to the telluric correction, we carry out separate analyses for , with the fixed number of SVD components removed from each of the orders, for each night. The motivation for the lower limit of = 3 is that, from this value onwards, the – maps of the combined transits (see Fig. 3) show clear detections of CO and HO near the expected planet position. The upper limit = 18 was chosen because this is where the SNR associated with the HO detection first drops below 4 (Fig. 4).
@@ 5.1 Four models of WASP-121b
To explore the effect of atmospheric drag (e.g., Showman et al. 2013; Komacek & Showman 2016; Parmentier & Crossfield 2018), we also consider two models with drag timescales s (weak drag) and s (strong drag), which are shown in the third and fourth column of Fig. 8. The drag timescale encapsulates a variety of physical mechanisms, such as turbulent mixing (Li & Goodman 2010), Lorentz-force braking of ionised winds in the planet’s magnetic field (Perna et al. 2010a), and Ohmic dissipation (Perna et al. 2010b). The drag deposits energy back into the atmosphere. As shown in Fig. 8, increasing the drag strength (and thus lowering ) slows down winds in the atmosphere131313 is a tunable parameter in the GCM. It is included as an additional “Rayleigh drag” term in the momentum equation solved by the dynamical core. This approach assumes that the drag force is uniform across the whole planet atmosphere. A drag-free scenario corresponds to .. This leads to less efficient heat redistribution and a more symmetric temperature structure. Also, the equatorial jet gets suppressed – in the strong-drag model, there is only a day-to-night flow.
Another notable difference between the model from Parmentier et al. (2018) and the models from Tan et al. (2024) is that the latter account for opacities from Fe when evaluating heating and cooling rates, making the thermal inversion on the dayside extend to much lower pressures. We refer to Table 1 in Tan et al. (2024) for the full list of opacities considered in their radiative transfer and to Freedman et al. (2014) for the opacities used by Parmentier et al. (2018). While the extra optical opacity causes the models with H dissociation/recombination to look substantially hotter, they have nearly the same effective temperature (2400 K) at the dayside photosphere (0.1 bar) as the model from Parmentier et al. (2018).
@@ 5.2 Computing absorption trails
Before feeding the GCM outputs into gCMCRT, we map the atmospheric structures onto a 3D grid with altitude (instead of pressure) as a vertical coordinate. We account for the fact that each atmospheric column has a different scale height, which is set by local gravity, temperature, and mean-molecular weight. For each of the four WASP-121b models we simulate 25 spectra (equidistant in orbital phase) between phase angles 17.6, which also cover the ingress and egress. We assume an edge-on orbit, a semi-major axis of 0.025 AU, a stellar radius of 1.46 , and an orbital period of 1.27 days. At each orbital phase angle, gCMCRT simulates a transmission spectrum by randomly shooting photon packets at the planet limb and evaluating the optical depth encountered by each photon packet. In this calculation, the code accounts for Doppler shifts imparted on the opacities by the radial component of the local wind vector and planet rotation (Wardenier et al. 2021). The transit depth at a given wavelength is then obtained by averaging over all photon packets. We use 105 photon packets per wavelength to accurately model the shapes, depths, and shifts of the spectral lines. Because we do not explicitly treat scattering, the propagation direction of the photon packets does not change throughout the calculation. As in Wardenier et al. (2023), the optical and infrared spectra are computed at resolutions and , respectively. For the radiative transfer, we include the same set of continuum opacities and line species as in Wardenier et al. (2023).
@@ 6.6 Non-detection of OH
Although we find a hint of OH in the atmosphere of WASP-121b (see Figs. 3, 14), the associated signal (SNR 2.5) is not significant. This is remarkable given that OH was confidently detected (SNR 6.1) in the atmosphere of WASP-76b based on one transit visit with CARMENES (; Landman et al. 2021), which is on a smaller, 3.5-m telescope. Since the equilibrium temperature of WASP-121b is 250 K higher than that of WASP-76b, one explanation could be that OH has further dissociated into atomic O and H on the hotter dayside of WASP-121b, such that its features in the transmission spectrum are weaker. A retrieval would allow to shed light on the upper limits of the OH abundance in the region probed by our transit data.
@@ Appendix A Supplementary figures
Figure 15 shows the results of the 2D VULCAN simulation described in Section 6.3. As demonstrated in the plot, chemical kinetics only has a strong impact on the water abundances in small regions around 10-5 bar, which is not enough to substantially impact the HO signal of the planet.